Seismic activity during this 24-hour period was similar to that recorded yesterday. Nine (9) earthquakes were located. These occurred south of St George's Hill (4 events at depths of 4-6km), beneath the Soufriere Hills (2 events; depths 2-3km), in the Windy Hill area (1 event; depth approximately 4km) and beneath the Centre Hills (1 event; 4km depth). There have been several very small long-period earthquakes recorded at the Gages seismometer. Some of these are reminiscent of the seismicity recorded during recent dome growth. These have been accompanied intermittently by larger long-period events. Broadband tremor has occurred almost continuously throughout the period.
Attempts were made to occupy the Long Ground-White's-Castle Peak and Farrell's-Windy Hill-St. George's Hill EDM triangles. However, very low cloud levels around the volcano prevented any measurements being made. The GPS sites at St George's Hill, Radio Antilles, Harris' Lookout and Chances Peak were occupied today, but the data are still being processed.
A helicopter flight late this afternoon allowed visual observation of the crater area. A new area of dome growth was identified in the moat area to the west of the July 18th vent. Reddish coloured boulders were observed to have spalled off the new dome and fallen to the north into the area formerly occupied by a pond of water. Vigorous steaming was observed from this site. An apron of light coloured material was observed to have fallen from a small ridge between the dome filling the July 18th vent and Castle Peak. Sulphur deposits were observed near steaming cracks on this feature. Spalling of material from the dome in the July 18th vent has also continued in a southerly direction.